RC Update and Run Plan for Aug 26, 2020

From HallCWiki
Jump to: navigation, search

Back to the precious page

RC Meeting Information

RC Daily Meetings are at 4:15 PM, remote

General updates

  • OK overnight running
    • Transport ion-chamber trip even without beam. Between 1am-3am , Hall B and D are up, we were down.
    • Current reduction: whenever Hall A is at 150uA, we could only take 24 uA max
    • 2L25 last four cavity stability issue
  • HMS dipole trip due to cryo ESR main return valve issue at 12pm today, became back around 2:30 today. SHMS kept taking data.
  • EPR calibration at longitudinal 180 is done
  • Moller completed yesterday
  • Ref-H2 Kin C-Z: completed! ~4.0 hours mostly at 30 uA.
  • Production started mid night. So far collected 6-7 hours of good data.
  • Current situation
    • Reference cell on H2
    • HMS is at 20 degrees -4 GeV and SHMS is at 18 degrees -5.6 GeV (Kin C-Z)
  • Last IHWP flip: 9:10am today
  • Last NMR: 15:33 today

Pending problem(s) / Things to track

We have two items that need to be addressed in the next opportunistic access (request already submitted to PD):

  • Resetting the SHMS rotation controller.
  • Diagnose/resolve the transport ion chamber issue. (Investigation is ongoing by Jerry Kowal)

Current reduction is also being worked on (seems to have improved after the MCC's initial attempt)

Monitoring

  • Compressor issue (spare compressor is on-site, but not connected yet) [Walter, Brad]
  • Spare reference cell options/status [Todd, JP]
  • Target mounting for Tommy (next cell) [Arun, et al]
  • Target mounting for Butterball [Arun, et al]
  • 1-pass changeover planning

Recent polarization measurements

@ 180 deg

  • Swing (08/25) 16:42 NMR - 39.5%
  • Swing (08/25) 23:50 NMR - 45.2%
  • Owl (08/26) 5:02 NMR - 45%
  • Day (08/26) 10:05 NMR - 45.8%, pNMR - 45.3%
  • Day (08/26) 15:33 NMR - 44.1%, pNMR - 46.2%
  • Two more pNMR measurements required

Current Run Plan

Normal Production running

  • Prerequisites
    • Make sure the raster size is correct (4--4.5mm)
    • Monitor beam parameters/configuration very often and post an entry like this at the beginning of your shift and when run conditions have changed
    • Running on cell targets: raster always ON (4--4.5mm, size will be determined by the scraping checks), orbit locks ON, current ramp rate 1 muA/sec
    • Always take 120Hz of EDTM
    • Fill the BTA with realistic numbers. If the desired beam current is not delivered prorate your ABUs!
    • Change the IHWP at least every 8 hours when running on glass cells. If Hall A hasn't changed in 8 hours, call the RC
  • Run plan
    • Completing Moller
    • NMR
    • If polarization > 40% Production
      • The beam position on the polarizated target is not yet verified
    • If polarization < 40% Continue to execute reference cell run plan
    • When switched back to production from pick-coil position always check with 5uA beam to check if the target cell clipping or not, before ask 30uA beam.

Production Run Plan on longitudinal (180 position) applies to Aug 26 day shift

Reference Cell Run Plan

 === REF-VAC on Kin-A and Kin-X (1--2 hours) ===  <Completed! 8/25 Owl> 
      HMS Kin-A:  p0 = -4.2 ; th=13.5
     SHMS Kin-X:  p0 = -7.5 ; th=11.0
   - Take 1--2 beam-on-target hours of data on the REF-VAC target
     - Use EL_CLEAN triggers in both arms (ps1=ps2=-1); Start with
       ps3=0, but adjust current and prescales to optimize as noted
       above.
 === REF-VAC on Kin-C and Kin-Z (3--4 hours) === <Completed! 8/25 Owl > 
      HMS Kin-C:  p0 = -4.0 ; th=20.0
     SHMS Kin-Z:  p0 = -5.6 ; th=18.0
   - Set the Spectrometers to the settings above
     - Follow the *updated* instructions in the How-Tos to rotate the
       Spectrometers
     - NOTE: Both sets of magnets can be directly ramped down to the
             new settings -- Do /not/ go to the MOL settings!
   - Target:  Reference cell with Vacuum (current setting)
   * Optimize the beam current and prescales as needed, but start with
     no prescales and evaluate the rates first.
     - MAXIMUM 30 uA (or as much as MCC can deliver and be stable)
   - Take 2 hours of beam-on-target data with
       * Use trigger T3 in *both* arms: set prescales
          ps3=0           (El_clean enabled, no prescaling)
          ps1=-1, ps2=-1  (disable both T1 and T2)
   - Take 2 hours of beam-on-target data with
       * Use trigger T1 in *both* arms: set prescales
          ps1=0           (3/4 enabled, no prescaling)
          ps2=-1, ps3=-1  (disable both T1 and T2)
 === REF-H2 on Kin-C and Kin-Z (3--4 hours) ===  <Completed! 8/25 Swing> 
      HMS Kin-C:  p0 = -4.0 ; th=20.0
     SHMS Kin-Z:  p0 = -5.6 ; th=18.0
   - Target:  Reference cell with 135psig of Hydrogen (H2)
   * Optimize the beam current and prescales as needed, but start with
     no prescales and evaluate the rates first.
     - MAXIMUM 30 uA (or as much as MCC can deliver and be stable)
   - Take 3-4 hours of beam-on-target data with
       * Use trigger T1 in *both* arms: set prescales
          ps1=0           (3/4 enabled, no prescaling)
          ps2=-1, ps3=-1  (disable both T1 and T2)
 === Optics on Kin-B and Kin-Z === <- We start here 
    (NOTE: You will NOT get here before the Moller -- that's OK.  This
           is really just notes for the shift *after* the Moller.)
      HMS Kin-C:  p0 = -4.0 ; th=20.0
     SHMS Kin-Z:  p0 = -5.6 ; th=18.0
       - Switch the target to C-OPTICS
       - Set the collimators to
            HMS: 'Sieve'
           SHMS: 'Centered Sieve'
       * Use trigger T3 in *both* arms: set prescales
          ps3=0           (El_clean enabled, no prescaling)
          ps1=-1, ps2=-1  (disable both T1 and T2)
       - Ask MCC:
         - to set the beam positions to 0,0
         - to set the raster to 0.5mm
         - request 40 uA (or as much as they can give us and keep
           reasonably stable beam)

Notices in Effect for Shift Crew

  • Follow all COVID rules and regulations as well as remind each other. Additional food guild lines:
    • All food are removed from the round table. Fruit are now store in the fridge, and individually packaged snacks are store on the shelf below the coffee machine.
    • No food near the SL and TO control station. If you would like to eat something, please use the empty round table (used to store food in the past), that is now empty and sanitized.
    • New coffee guide line: it is much preferred if you can bring you mug for the coffee. During the shift if one need to drink coffee near the control station, then the coffee mug must have a lid or have some kind of cover.
  • Beam positions: [1]
  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[2]
  • An new reference Cell refill procedure: New procedure
  • Magnet policies
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • Slow ramping procedure of the SHMS, go small steps to the MOL, an great example (1 hour 30 min process)
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
    • New updates to the magnet momentum ramping procedure: https://hallcweb.jlab.org/wiki/index.php/How_to_operate_and_monitor_the_Spectrometer_Magnets
    • If any of the Spectrometers magnets trip, take a screenshot of the PSU window and check the hall for smoke or water with the camera before resetting the interlock and post it to hclog.
  • New SHMS matrix element
  • When rotating the spectrometer, use HCCAM01 to keep an eye on the SOLID Cherenkov test stand when rotating SHMS (in addition to watching the live angle-cam image, etc). The updated rotation instruction can be found: https://hallcweb.jlab.org/wiki/index.php/Spectrometer_Rotation#NOTICES_IN_EFFECT

Planned Accelerator Configuration Changes

  • Hall C to 1st pass on 8/31 and back to 5th pass on 9/03


  • Bluejeans remote connection information:
 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920